# Overparameterized Multiple Linear Regression as Hyper-Curve FittingE. Atza and N. Budko

### v1

@@ Abstract
The paper shows that the application of the fixed-effect multiple linear regression model to an overparameterized dataset is equivalent to fitting the data with a hyper-curve parameterized by a single scalar parameter. This equivalence allows for a predictor-focused approach, where each predictor is described by a function of the chosen parameter. It is proven that a linear model will produce exact predictions even in the presence of nonlinear dependencies that violate the model assumptions. Parameterization in terms of the dependent variable and the monomial basis in the predictor function space are applied here to both synthetic and experimental data. The hyper-curve approach is especially suited for the regularization of problems with noise in predictor variables and can be used to remove noisy and ‘improper’ predictors from the model.
@@
overparametrization, linear regression, model selection, variable selection, regularization.
{MSCcodes}
15A06, 62R07, 68T01, 65F22, 62J20.
@@ 1 Introduction
Due to its column-centered nature, the hyper-curve approach allows to filter out the predictors (features) that are either too noisy or do not satisfy the topological requirements of a linear model, which significantly improves the predictive power of the trained linear model, removes the features that may otherwise introduce the illusion of understanding [21], and suggests the subsets of predictors where a non-linear or a higher-dimensional-manifold model would be more adequate.
The paper is structured as follows. In Section 2 we define Fundamentally Overparameterized Datasets, MLR, PARametric hyper-CURve (PARCUR) and IR models, prove their equivalence and identify the condition for the exact prediction of a test dataset. In Section 3 we study the behavior of the polynomial IR model applied to a dataset which contains both the polynomial predictors as well as predictors that have a non-functional relation to the dependent variable. We establish conditions under which such a dataset is a FOP dataset. In Section 4 we consider noisy data and introduce a polynomial degree truncation regularization scheme that can handle the noise in both the dependent and predictor variables. In Section 5 we propose a novel predictor removal algorithm that does not suffer from the ambiguities common to heuristic feature selection methods. In Section 6 we apply the regularized IR model with predictor removal to the widely available experimental chemometric Yarn dataset [22], [23], and demonstrate the presence of both curve-like and higher-dimensional manifolds in this dataset. Finally, we present our conclusions and discuss the possible extensions of the PARCUR and IR models.
@@ 2 Parametric Hyper-Curve and Inverse Regression models
In particular, the Multiple Linear Regression (MLR) model,
@@
Our main results, expressed in Theorem 2.3 and Theorem 2.5, are the conditions on the exact prediction by the MLR model and the equivalence of the predictions made by the MLR and PARCUR models for what we call a Fundamentally OverParameterized (FOP) dataset. This equivalence allows to analyze the different types of column functions $x_{j}(s)$, $j=1,\dots,p$, and establish the conditions on the existence of a FOP dataset for different types of predictor data.
In practice, any regression model is trained on a finite discrete dataset. An overparameterized dataset arises whenever the number $n$ of training samples is smaller than the number $p$ of parameters or predictors. This can happen simply due to the lack of experimental data and additional data can transform an overparameterized dataset into a well-defined or even an underparameterized one. However, in the present paper, we shall focus on the more fundamental case, where the simple addition of the training data does not change the overparameterized nature of the dataset.
@@
Notice, we use the same notation for the predictor data $x_{i,j}$ and the column functions $x_{j}(s)$, with the relation being $x_{i,j}=x_{j}(s_{i})$. Also, obviously, $s=y$ in the IR model (4).
@@
where the elements $[A]_{k,j}=a_{k,j}$ of the matrix $A\in{\mathbb{R}}^{n\times p}$ and the elements $[{\bm{a}}_{0}]_{k}=a_{k,0}$ of the vector ${\bm{a}}_{0}\in{\mathbb{R}}^{n}$ are the expansion coefficients from (9).
Formally, the PARCUR model (4) can be trained by computing the matrix $A$ and the vector ${\bm{a}}_{0}$ as:
@@
If the training set ${S}_{q}$ is not complete, then there exists a data row ${\bm{s}}_{\rm t}\in{\mathcal{S}}$ that is not a linear combination of the rows of $S_{q}$.
@@
Proof 2.6.
The proof follows from substituting the relations (15) into (18).
If the training dataset is incomplete and the testing data-matrix $S_{\rm t}$ is not (yet) in the row-range of the training data-matrix $S_{n}$, there will be an error in the predictions made by the MLR and PARCUR models. Let $S_{\rm t}$ be decomposed as
@@ 3 Data containing polynomial column functions
This also puts the PARCUR model into the context of polynomial fitting. Whether the columns of $X$ have or have not been generated by polynomial functions of $y$, the IR model with the basis matrix (28) will be projecting all columns of $X$ on the monomial basis. From the computational point of view, Vandermonde matrices, while theoretically invertible, are hard to work with for sizes above $n=15$ and become the source of significant round-off errors. Luckily, one normally does not need polynomial functions of very high degree to adequately describe a data column. To minimize the numerical errors, we also normalize the range of the $y$-data to fit within the interval $[-1,1]$.
@@
    $x_{j}(y)$ is a polynomial in $y$ of degree less or equal to $n-1$
@@
    $x_{j}(y)$ is a polynomial in $y$ of degree higher than $n-1$
@@
A non-functional dependence between the predictor $x_{j}$ and the dependent variable $y$ would emerge if the data was situated on a curve that cannot be parameterized by $y$, Fig. 1 (left). Choosing a different parameterization could transform these ’nonfunctional’ data to functions $y(s)$, $x_{j}(s)$, $j=1,\dots,p$, as in the general PARCUR formulation. A more severe case of non-functional dependence arises where the data is situated on a higher-dimensional manifold, such as a hyper-surface, Fig. 1 (right), and no alternative parameterization can fix this problem. The column data that one observes in such ‘nonfunctional’ cases are illustrated in Fig. 1 (bottom, two-dimensional scatter plots).
@@
Let ${\mathcal{S}}$ be a dataset with one independent variable $y$ and $p$ predictor variables $x_{j}$, $j=1,\dots,p$. Let also ${\mathcal{S}}$ contain $k\leq p$ predictors that are polynomials in $y$ of degrees greater than $q-1$, or have a non-functional relation to $y$. Then, ${\mathcal{S}}$ is a fundamentally overparameterized dataset (in linear sense) of rank $q$, $2\leq q\leq p$, if its remaining $p-k$ predictors are the polynomials in $y$ of degree $r$, $11\leq r\leq q-k-$ 1.
Proof 3.2.
@@
where $V^{-1}{\bm{y}}={\bm{e}}_{2}\in{\mathbb{R}}^{m}$ is the second standard basis vector and $O\in{\mathbb{R}}^{(m-q+k)\times(p-k)}$ is the matrix of all zeros. Here we have used the fact that the polynomial fit to non-functional data may produce a polynomial of degree greater than $q-1$. It is obvious that ${\rm rank}(A_{1,1})\leq(q-k)$ and, for any $m\geq(q-k)$, ${\rm rank}(A_{2,2})\leq k$. Therefore, ${\rm rank}(A)\leq q$, and ${\rm rank}(S_{m})\leq q$ for any $m$, showing that ${\mathcal{S}}$ is a FOP dataset.
The above Theorem 3.1 shows that having extremely high-degree polynomials, e.g., with degrees higher than $p-1$ and non-functional dependencies among the predictors does not prevent the MLR and the IR models from making exact predictions as long as there are also polynomial data of sufficiently low degree and the set on which the model is trained is complete. Moreover, a complete dataset can, in principle, be achieved with $n<p$ samples. The latter fact may seem surprising as it appears that we are able to recover a polynomial of degree higher than $n-1$ or a non-functional dependence by training on just $n$ data points. However, it becomes less surprising if we consider the form of the MLR and IR predictors given by (6) and (18), as in both cases the leftmost matrix $X_{\rm t}$ contains the $X$-data from the test dataset which ones is trying to ‘predict’.
@@
Fig. 2 (top) illustrates the performance of the IR model with the three classes of column data discussed above. Specifically, we are considering the cases where: all predictors are polynomial functions in $y$ of degree $r-1\leq q$ (blue, circles), some of the predictors are high-degree polynomials in $y$ (orange, squares), and some of the predictors have non-functional relation to the dependent variable $y$ (green, triangles).
Figure 2: Top: prediction errors of the IR model as functions of the training dataset rank obtained with invertible basis matrix $V$ and exact data (see text for full explanation). Bottom: test of the regularization algorithm on exact (noiseless) data and over-complete training dataset. Bright enlarged colored markers correspond to the minima of the validation errors that indicate the optimal polynomial degree $r^{*}$ along the horizontal axis (bottom, right) and show the values of the test errors attained with these $r^{*}$ (bottom, left). Note, the horizontal axis displays ${\rm rank}(\hat{A})=r+1$, rather than the actual polynomial degree $r$.
In these numerical experiments the data $x_{j}(y)$ are generated by randomly sampling the range of $y\in[-1,1]$ and evaluating polynomial functions of various degrees at the sampled points. Columns that are not functions of $y$ are generated as non-invertible functions $y(x_{j})$, similar to those in the examples of Fig. 1. The predictions are computed as in Eq.’s (18), where the coefficient matrix $A$ is obtained from the training $X$-data as in Eq. (15).
@@
Fig. 2 shows the prediction errors as functions of the training set rank (top row), and of the polynomial representation degree (bottom row). The bright colored dashed lines give the training set errors and the dim solid gray lines show the corresponding errors on the test dataset. Plots on the left show the dependent variable errors and on the right – the errors for the predictor variables. As expected, the errors drop as soon as the training dataset becomes complete. With the high-degree polynomial predictors and especially with non-functional predictors the errors on the training set begin to rise at the end and the errors on the test set do not drop to machine precision values as happens with the low-degree polynomial predictors. This is due to the fact that the coefficient matrices in the former two cases become ill-conditioned, making the predictors more sensitive to numerical round-off errors.
@@ 4 Polynomial regularization
Fig. 2 (bottom) illustrates the testing of the polynomial-degree truncation regularization scheme for the IR model on the three noiseless datasets described in Section 3. The meaning of the lines and symbols is the same as in Fig. 2 (top), see Section 3. The errors $\rho({\bm{y}}_{\rm v})$ (left, colored, dashed) and $\rho(X_{\rm v})$ (right, colored, dashed) on the training dataset are the mean CV errors (36). All data are exact up to numerical precision. The only differences with the IR model of Fig. 2 (top) are the over-complete nature of the training datasets ($n=150$) and the application of the LS estimates (23) of the coefficient matrix $\hat{A}_{r}$. Therefore, the horizontal axis in Fig. 2 is the number $r$ of columns in the rectangular basis matrix $V\in{\mathbb{R}}^{n\times r}$, which is no longer equal to the number of samples $n$.
In the tests of Fig. 2 (bottom-left), the $\rho({\bm{y}}_{\rm v})$ error starts at the level of machine precision for $r=2$ and increases thereafter due to the accumulation of numerical errors. That the $\rho({\bm{y}}_{\rm v})$ error grows for $r>2$ has to do with the fact that the exact representation of ${\bm{y}}$ is attained already at $r=2$ in the monomial basis. The enlarged coloured markers in the right plot indicate the minima of the $\rho(X_{\rm v})$, the same markers in the left plot show the levels of $\rho({\bm{y}}_{\rm t})$ errors attained with the corresponding $r^{*}$. As can be seen, the optimal $r^{*}$ gives the smallest achievable error $\rho({\bm{y}}_{t})$. Hence, in this noiseless case the regularization procedure correctly identifies the rank of the complete dataset as the optimal $r^{*}$.
@@
Fig. 3 (top) illustrates the application of the regularization procedure to the problem where only the dependent-variable data $\bm{y}$ contains additive Gaussian noise. The columns of the matrix $X$ are exact up to machine precision. We consider three different standard deviations $\sigma=0.05,0.1$, and $0.2$, corresponding to $5\%$, $10\%$, and $20\%$ levels of noise relative to the ${\bm{y}}$-data magnitude. As can be seen from the test-error curves (top-left plot, dim solid gray), the errors $\rho({\bm{y}}_{\rm t})$ attained with these choices of $r^{*}$ (top-right plot, enlarged colored squares) are close to the smallest achievable $\rho({\bm{y}}_{\rm t})$ errors for all considered noise levels. At the same time, the minimal $\rho({X}_{\rm v})$ errors do not correspond to the smallest achievable $\rho({X}_{\rm t})$ errors (top-right plot, dim solid gray curves).
Fig. 3 (bottom) illustrates the case of additive noise in the dependent variable and in one-third of the predictors. While the behavior is generally similar to the previous case, the attained optimal $\rho({\bm{y}}_{\rm t})$ errors increase more rapidly with the level of noise. The growth of the $\rho({\bm{y}}_{\rm v})$ error (bottom-left, colored dashed) is now caused not only by the accumulation of numerical errors and the statistical noise in ${\bm{y}}$ (via $V$), but by the statistical noise in the $X$-data as well (via $\hat{A}$).
Figure 3: Top: application of the regularization procedure in the case of exact predictor data $X$ and noisy dependent-variable data ${\bm{y}}+{\bm{\epsilon}}$, ${\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},\sigma_{i}^{2}I)$, with $\sigma_{1}=0.05$ (blue), $\sigma_{2}=0.1$ (orange), and $\sigma_{3}=0.2$ (green). Bright colored lines: validation errors $\rho({\bm{y}}_{\rm v})$ (left, right vertical axis) and $\rho(X_{\rm v})$ (right). Dim dashed gray lines: errors on the test dataset, $\rho({\bm{y}}_{\rm t})$ (left, left vertical axis) and $\rho(X_{\rm t})$ (right). Bright colored square markers: values of the test errors $\rho({\bm{y}}_{\rm t})$ and $\rho({X}_{\rm t})$ attained at $r^{*}$ (left and right respectively). Bright colored diamond markers: values of the test errors $\rho({\bm{y}}_{\rm t})$ (left, left vertical axis) and $\rho({X}_{\rm t})$ (right) attained with $r^{*}$ and the optimal set $\{p\}_{\text{opt}}$ of predictors. Bottom: same as in the row above, but with a third of the columns of $X$ affected by the additive Gaussian noise of the same type as the noise in the dependent variable ${\bm{y}}$.

### v2

@@ Abstract
This work demonstrates that applying a fixed-effect multiple linear regression (MLR) model to an overparameterized dataset is mathematically equivalent to fitting a hyper-curve parameterized by a single scalar. This reformulation shifts the focus from global coefficients to individual predictors, allowing each to be modeled as a function of a common parameter. We prove that this overparameterized linear framework can yield exact predictions even when the underlying data contains nonlinear dependencies that violate classical linear assumptions. By employing parameterization in terms of the dependent variable and a monomial basis, we validate this approach on both synthetic and experimental datasets. Our results show that the hyper-curve perspective provides a robust framework for regularizing problems with noisy predictors and offers a systematic method for identifying and removing ’improper’ predictors that degrade model generalizability.
@@
hyper-curve fitting, high-dimensional regression, dimensionality reduction, model selection, variable selection, basis-function expansion, implicit regularization, generalization error
articletype: Paper
@@ 1 Introduction
Due to its variable-centered nature, the hyper-curve approach allows to filter out the predictors (features) that are either too noisy or do not satisfy the topological requirements of a linear model, which significantly improves the predictive power of the trained linear model, removes the features that may otherwise introduce the illusion of understanding [12], and suggests the subsets of predictors where a non-linear or a higher-dimensional-manifold model would be more appropriate.
The paper is structured as follows. In Section 2 we define Fundamentally OverParameterized (FOP) datasets, MLR, PARametric hyper-CURve (PARCUR) and IR models, prove their equivalence and identify the condition for the exact prediction of a test dataset. In Section 3 we study the behavior of the polynomial IR model applied to a dataset which contains both the polynomial predictors as well as predictors that have a non-functional relation to the dependent variable. We establish conditions under which such a dataset is a FOP dataset. In Section 4 we consider noisy data and introduce a polynomial degree truncation regularization scheme that can handle the noise in both the dependent and predictor variables. In Section 5 we propose a novel predictor removal algorithm that does not suffer from the ambiguities common to heuristic feature selection methods. In Section 6 we apply the regularized IR model with predictor removal to the widely available experimental chemometric Yarn dataset [13], [18], and demonstrate the presence of both curve-like and higher-dimensional manifolds in this dataset. Finally, we present our conclusions and discuss the possible extensions of the PARCUR and IR models.
@@ 2 Parametric Hyper-Curve and Inverse Regression models
In particular, the MLR model,
@@
Our main results, expressed in theorem 2.1 and theorem 2.2, are the conditions on the exact prediction by the MLR model and the equivalence of the predictions made by the MLR and PARCUR models for what we call a FOP dataset. This equivalence allows to analyze the different types of column functions $x_{j}(s)$, $j=1,\dots,p$, and establish the conditions on the existence of a FOP dataset for different types of predictor data.
In practice, any regression model is trained on a finite discrete dataset. An overparameterized dataset arises whenever the number $n$ of training samples is smaller than the number $p$ of parameters or predictors. This can happen simply due to the lack of experimental data and additional experiments could transform an incidentally overparameterized dataset into a well-defined or even an underparameterized one. However, in the present paper, we shall focus on the more fundamental case, where the simple addition of the training data does not change the overparameterized nature of the dataset. To do so, we imagine that our dataset is potentially infinite, i.e., one can draw or generate data from this dataset as many times as one wishes, e.g., by simply performing the measurements of the same physical quantities over and over again.
@@
Notice that we use the same notation for the predictor data $x_{i,j}$ and the column functions $x_{j}(s)$, with the relation being $x_{i,j}=x_{j}(s_{i})$. Also, obviously, $s=y$ in the IR model (4).
@@
where the elements $A]_{k,j}=a_{k,j}$ of the matrix $A\in{\mathbb{R}}^{n\times p}$ and the elements ${\bm{a}}_{0}]_{k}=a_{k,0}$ of the vector ${\bm{a}}_{0}\in{\mathbb{R}}^{n}$ are the expansion coefficients from (9).
Formally, the PARCUR models (3) and (4) can be trained by computing the matrix $A$ and the vector ${\bm{a}}_{0}$ as:
@@
If the training set ${S}_{q}$ is not complete, then there exists a data row ${\bm{s}}_{\rm t}\in{\mathcal{S}}$ that is not a linear combination of the rows of $S_{q}$. ∎
@@
Proof.
The proof follows from substituting the relations (12) into (15). ∎
If the training dataset is incomplete and the rows of the testing data-matrix $S_{\rm t}$ are not (yet) in the span of the rows of the training data-matrix $S_{n}$, there will be an error in the predictions made by the MLR and PARCUR models. Let $S_{\rm t}$ be decomposed as
@@ 3 Data containing polynomial column functions
This also puts the PARCUR model into the context of polynomial fitting. Whether the columns of $X$ have or have not been generated by polynomial functions of $y$, the IR model with the basis matrix (21) will be projecting all columns of $X$ on the monomial basis. From the computational point of view, Vandermonde matrices, while theoretically invertible, are hard to work with for sizes above $n=15$ and become the source of significant round-off errors. Luckily, one normally does not need polynomial functions of very high degree to adequately describe a data column. To minimize the numerical errors, we also normalize the range of the $y$-data to fit within the interval [1,1][-1,1].
@@
    $x_{j}(y)$ is a polynomial in $y$ of degree less or equal to n1n-1
@@
    $x_{j}(y)$ is a polynomial in $y$ of degree higher than n1n-1
@@
A non-functional dependence between the predictor $x_{j}$ and the dependent variable $y$ would emerge if the data was situated on a curve that cannot be parameterized by $y$, Figure 1 (left). Choosing a different parameterization could transform these ’nonfunctional’ data to functions $y(s)$, $x_{j}(s)$, $j=1,\dots,p$, as in the general PARCUR formulation. A more severe case of non-functional dependence arises where the data is situated on a higher-dimensional manifold, such as a hyper-surface, Figure 1 (right), and no alternative parameterization can fix this problem. The column data that one observes in such ‘nonfunctional’ cases are illustrated in Figure 1 (bottom, two-dimensional scatter plots).
@@
Let $\mathcal{S}$} be a dataset with one independent variable $y$ and $p$ predictor variables $x_{j}$, $j=1,\dots,p$. Let also $\mathcal{S}$} contain $k\leq p$ predictors that are polynomials in $y$ of degrees greater than q1q-1, or have a non-functional relation to $y$. Then, $\mathcal{S}$} is a fundamentally overparameterized dataset (in linear sense) of rank $q$, $2\leq q\leq p$, if its remaining pkp-k predictors are the polynomials in $y$ of degree $r$, $1\leq r\leq q-k-1$.
Proof.
@@
where $V^{-1}{\bm{y}}={\bm{e}}_{2}\in{\mathbb{R}}^{m}$ is the second standard basis vector and $O\in{\mathbb{R}}^{(m-q+k)\times(p-k)}$ is the matrix of all zeros. Here we have used the fact that the polynomial fit to non-functional data may produce a polynomial of degree greater than q1q-1. It is obvious that ${\rm rank}(A_{1,1})\leq(q-k)$ and, for any $m\geq(q-k)$, ${\rm rank}(A_{2,2})\leq k$. Therefore, ${\rm rank}(A)\leq q$, and ${\rm rank}(S_{m})\leq q$ for any $m$, showing that $\mathcal{S}$} is a FOP dataset. ∎
The above Theorem 3.1 shows that having extremely high-degree polynomials, e.g., with degrees higher than p1p-1 and non-functional dependencies among the predictors does not prevent the MLR and the IR models from making exact predictions as long as there are also polynomial data of sufficiently low degree and the set on which the model is trained is complete. Moreover, a complete dataset can, in principle, be achieved with $n<p$ samples. The latter fact may seem surprising as it appears that we are able to recover a polynomial of degree higher than n1n-1 or a non-functional dependence by training on just $n$ data points. However, it becomes less surprising if we consider the form of the MLR and IR predictors given by (6) and (15), as in both cases the leftmost matrix $X_{\rm t}$ contains the $X$-data from the test dataset which ones is trying to ‘predict’.
@@
Figure 2 (top) illustrates the performance of the IR model with the three classes of column data discussed above. Specifically, we are considering the cases where: all predictors are polynomial functions in $y$ of degree $r-1\leq q$ (blue, circles), some of the predictors are high-degree polynomials in $y$ (orange, squares), and some of the predictors have non-functional relation to the dependent variable $y$ (green, triangles).
Figure 2: Drop in prediction errors of IR models when the training set becomes complete (noiseless data, left – dependent variable $\bm{y}$}, right – independent predictor variables $X$). Colored lines – training set errors, gray lines – test set errors. Blue, circles: maximum polynomial degree of predictors is $11$. Orange, squares: maximum polynomial degree of predictors is $14$. Green, triangles: some of the predictors have non-functional dependence on the dependent variable. Top: square, invertible basis matrix $V$. Bottom: non-square basis matrix $V$ (over-complete training dataset) and a formal test of the polynomial truncation regularization algorithm with noiseless data. The identified optimal polynomial degrees $r^{*}$ coincide with the ranks of the corresponding complete training datasets (bright enlarged colored markers show the minima of validation errors).
In these numerical experiments the data $x_{j}(y)$ are generated by randomly sampling the range of $y\in[-1,1]$ and evaluating polynomial functions of various degrees at the sampled points. Columns that are not functions of $y$ are generated as non-invertible functions $y(x_{j})$, similar to those in the examples of Figure 1. The predictions are computed as in Eq.’s (15), where the coefficient matrix $A$ is obtained from the training $X$-data as in Eq. (12).
@@
Figure 2 shows the prediction errors as functions of the training set rank (top row), and of the polynomial representation degree (bottom row). The bright colored dashed lines give the training set errors and the dim solid gray lines show the corresponding errors on the test dataset. Plots on the left show the dependent variable errors and on the right – the errors for the predictor variables. As expected, the errors drop as soon as the training dataset becomes complete. With the high-degree polynomial predictors and especially with non-functional predictors the errors on the training set begin to rise at the end and the errors on the test set do not drop to machine precision values as happens with the low-degree polynomial predictors. This is due to the fact that the coefficient matrices in the former two cases become ill-conditioned, making the predictors more sensitive to numerical round-off errors.
@@ 4 Polynomial regularization
Figure 2 (bottom) illustrates the testing of the polynomial-degree truncation regularization scheme for the IR model on the three noiseless datasets described in Section 3. The meaning of the lines and symbols is the same as in Figure 2 (top), see Section 3. The errors $\rho({\bm{y}}_{\rm v})$ (left, colored, dashed) and $\rho(X_{\rm v})$ (right, colored, dashed) on the training dataset are the mean CV errors (27). All data are exact up to numerical precision. The only differences with the IR model of fig. 2 (top) are the over-complete nature of the training datasets ($n=150$) and the application of the LS estimates (20) of the coefficient matrix $\hat{A}_{r}$. Therefore, the horizontal axis in Figure 2 is the number $r$ of columns in the rectangular basis matrix $V\in{\mathbb{R}}^{n\times r}$, which is no longer equal to the number of samples $n$.
In the tests of Figure 2 (bottom-left), the $\rho({\bm{y}}_{\rm v})$ error starts at the level of machine precision for $r=2$ and increases thereafter due to the accumulation of numerical errors. That the $\rho({\bm{y}}_{\rm v})$ error grows for $r>2$ has to do with the fact that the exact representation of $\bm{y}$} is attained already at $r=2$ in the monomial basis. The enlarged coloured markers in the right plot indicate the minima of the $\rho(X_{\rm v})$, the same markers in the left plot show the levels of $\rho({\bm{y}}_{\rm t})$ errors attained with the corresponding $r^{*}$. As can be seen, the optimal $r^{*}$ gives the smallest achievable error $\rho({\bm{y}}_{t})$. Hence, in this noiseless case the regularization procedure correctly identifies the rank of the complete dataset as the optimal $r^{*}$.
@@
Figure 3 illustrates the application of the regularization procedure to the problem where only the dependent-variable data $\bm{y}$ contains additive Gaussian noise. The columns of the matrix $X$ are exact up to machine precision. We consider three different standard deviations $\sigma=0.05,0.1$, and $0.2$, corresponding to $5\%$, $10\%$, and $20\%$ levels of noise relative to the $\bm{y}$}-data magnitude. As can be seen from the test-error curves (top-left plot, dim solid gray), the errors $\rho({\bm{y}}_{\rm t})$ attained with these choices of $r^{*}$ (top-right plot, enlarged colored squares) are close to the smallest achievable $\rho({\bm{y}}_{\rm t})$ errors for all considered noise levels. At the same time, the minimal $\rho({X}_{\rm v})$ errors do not correspond to the smallest achievable $\rho({X}_{\rm t})$ errors (top-right plot, dim solid gray curves).
Figure 4 illustrates the case of additive noise in the dependent variable and in one-third of the predictors. While the behavior is generally similar to the previous case, the attained optimal $\rho({\bm{y}}_{\rm t})$ errors increase more rapidly with the level of noise. The growth of the $\rho({\bm{y}}_{\rm v})$ error (bottom-left, colored dashed) is now caused not only by the accumulation of numerical errors and the statistical noise in $\bm{y}$} (via $V$), but by the statistical noise in the $X$-data as well (via $\hat{A}$).
Figure 3: Regularization and improper-variable removal in the case of noisy dependent-variable data ${\bm{y}}+{\bm{\epsilon}}$, ${\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},\sigma_{i}^{2}I)$, with $\sigma_{1}=0.05$ (blue), $\sigma_{2}=0.1$ (orange), and $\sigma_{3}=0.2$ (green). Predictor data $X$ is noiseless. Bright dashed colored lines: validation errors $\rho({\bm{y}}_{\rm v})$ (left plot, right vertical axis) and $\rho(X_{\rm v})$ (right plot). Dim solid gray lines: errors on the test dataset, $\rho({\bm{y}}_{\rm t})$ (left plot, left vertical axis) and $\rho(X_{\rm t})$ (right plot). Bright colored square markers: values of the test errors $\rho({\bm{y}}_{\rm t})$ and $\rho({X}_{\rm t})$ attained at $r^{*}$. Bright colored diamond markers: values of the test errors attained with $r^{*}$ and the optimal set $\{p\}_{\text{opt}}$ of predictors (improper variables removed).
Figure 4: Regularization and improper-variable removal with additive Gaussian noise present in both the dependent-variable $\bm{y}$} as well as in one third of predictor variables (columns of $X$). Levels of noise and plot legend are the same as in Figure 3